On March 1, 2019, Myriad Entertainment purchased a parcel of land for $950,000 by paying $490,000 in cash and signing a short-term note payable with the seller for $460,000. You must repay the $460,000 in exactly one year on March 1, 2020. You agree to pay the seller 5% interest (annual rate) on a quarterly basis (June 1, September 1, December 1, 2019; and March 1, 2020). Give the journal entries for each payment and the adjusting entry needed on December 31, 2019.
Answer #1

Answer:

Given information:

Land purchased=$9,50,000(cash basis=4,90,000 & note basis=4,60,000).

Interest 5% on quarterly basis calculated

The following are the journal entries of required payments:

Date Particulars Debit Credit
Mar 1 2019 Land $9,50,000
Cash $4,90,000
Note payable $4,60,000
(Narration: land purchased on both cash and credit basis)
June 1 2019 Interest expense(4,60,000*5/100*3/12=5,750) $5,750
Interest payable $5,750
(Narration:Interest payable on quarterly basis )
sep 1 2019 Interest expense(4,60,000*5/100*3/12=5,750) $5,750
Interest payable $5,750
(Narration:Interest payable on quarterly basis )
Dec 1 2019 Interest expense(4,60,000*5/100*3/12=5,750) $5,750
Interest payable $5,750
(Narration:Interest payable on quarterly basis )
ADJUSTING ENTRY FOR DEC 31 2019
Dec 31 2019 Interest expense(4,60,000*5/100*1/12=1917) $1,917
interest payable $1,917
(Narration:interest payable for one month dec)
Mar 1 2020 Interest expense(4,60,000*5/100*2/12=3,833) $3,833
Interest payable(5,750+5,750+5,750+1,917=19,167) $19,167
Note payable $4,60,000
Cash $4,83,000
(Narration:total note amount paid including total interest)